OpenStudy (anonymous):

Jim had a goal of riding his bike for an average (mean) of 90 miles each week. The number of miles he rode each of the first four weeks were 87, 81,92, and 95. Which value is the minimum number of miles he must ride during the fifth week to meet his goal? A. 89 B. 91 C. 95 D. 102

OpenStudy (anonymous):

I am not sure but I think it is the letter C. ?

OpenStudy (texaschic101):

(87 + 81 + 92 + 95 + x) / 5 = 90 (355 + x) / 5 = 90 --- multiply by 5 355 + x = 450 x = 450 - 355 x = 95 you are correct..it is C

OpenStudy (anonymous):

Thank you!

OpenStudy (texaschic101):

no problem :)